Determine the magnitude and directions of the currents in each resistor4) Determine the magnitudes and directions oft R1 25 + o 24001 9.0 V T- W R 68 R3 35 Q To:: 12V 0 0 05 GIT c 1S 0.
Answer #1

7 Ri=251 mizda R2=68s of I 3 § Rz=354 t I ~ Initially Assume current directions like figure shown. ► Al node , apply kischoff> 93 Jg = 9+9553 Iz=( +2513] - © in 3533+ 68[ 9177573 = 12 3513 + 6.58 +18.281; = 12 53.28112–6.58 Iz = 5.43 53.28 magnitude